# Benchmarking Predictive Coding Networks – Made Simple

### v1

@@ 4.1 Discriminative Mode
Here, we test the performance of PCNs on image classification tasks. We compare PC against BP, using both Squared Error (SE) and Cross Entropy (CE) loss, by adapting the energy function as described in (Pinchetti et al., 2022). For the experiments on MNIST and FashionMNIST, we use feedforward models with $3$ hidden layers of $128$ hidden neurons, while for CIFAR10/100 and Tiny ImageNET, we compare VGG-like models. We performed a large hyperparameter search over learning rates, optimizers, activation functions, and algorithm-specific parameters. All the details needed to reproduce the experiments, as well as a discussion about ‘lessons learned’ during such a large search, are in the Appendix B. Results, averaged over $5$ seeds are reported in Tab. 1.
Discussion.
The results show that the best performing algorithms, at least on the most complex tasks, are the ones where the target is nudged towards the real label, that are PN and NN. This is in line with previous findings in the Eqprop literature (Scellier et al., 2024). The recently proposed iPC, on the other hand, performs well on small architectures, as it is the best performing one on MNIST and FashionMNIST, but its performance worsten when it comes to train on large architectures. More broadly, the performance are comparable to those of backprop, except on the largest model. An interesting observation, is that all the best results for PC have been achieved using a VGG5, that has always outperformed the deeper VGG7. Future work should investigate the reason of such a phenomenon, as scaling up to more complex datasets will require the use of much deeper architectures, such as ResNets (He et al., 2016). In Section 5, we analyze possible causes, as well as comparing the wall-clock time of the different algorithms.
@@ Discussion.
Figure 2: CIFAR10 image reconstruction via autoencoding convolutional networks. In order: original, PC, iPC, BP, BP (half of the parameters).
Figure 3: Generative samples obtained by MCPC. Left: Contour plot of learned generative distribution compared to Iris data samples (x). Right: Samples obtained for a PCN trained on MNIST. In order: unconditional generation, conditional generation (odd), conditional generation (even). For more samples, please refer to Appendix C.2.
@@ 4.2 Generative Mode
We test the performance of PCNs on image generation tasks. We perform three different kinds of experiments: (1) generation from a posterior distribution; (2) generation via sampling from the learned joint distribution; and (3) associative memory retrieval. In the first case, we provide a test image $y$ to a trained model, run inference to compute a compressed representation $\bar{x}$ (stored in the latent vector $h_{0}$ at convergence), and produce a reconstructed $\bar{y}=h_{L}$ by performing a forward pass with $h_{0}=\bar{x}$). The model we consider are three layer networks. As this is an autoencoding task, we compare against autoencoders with three layer encoder/decoder structure (so, six layers in total). In the case of MNIST and FashionMNSIT, we use feedforward layers, in the case of CIFAR10 and CelebA, deconvolutional (and convolutional for the encoder) ones. The results in Tab. 2 and Fig. 2 report comparable performance, with a small advantage for PC compared to BP on the more complex convolutional tasks. In this case, iPC is the best performing algorithm, probably due to the small size of the considered models which allows for better stability. Furthermore, note that the BP architectures have double the amount of parameters, being the PC networks decoder only. If we halve the number of features in each layer of the autoencoder architecture (while keeping the bottleneck dimension unchanged), we get significantly reduced performance for BP (Fig. 2, bottom), achieving a loss of $10.66^{\pm{0.94}}\times 10^{-3}$ on CIFAR10. Details about these and the following experiments are provided in Appendix C.
Table 3: MSE (×$\times 10^{-4}$) of associative memory tasks. Columns indicate the number of hidden neurons while rows shows the training images to memorize. Results over 5 seeds.
@@
For the second category of experiments, we tested the capability of PCNs to learn, and sample from, a complex probability distribution. MCPC extends PC by incorporating Gaussian noise to the activity updates of each neuron. This change enables a PCN to learn and generate samples analogous to a variational autoencoders (VAE). This change shifts the inference of PCNs from a variational approximation to Monte Carlo sampling of the posterior using Langevin dynamics. Data samples can be generated from the learned joint $P_{\theta}(h)$ by leaving all states $h_{l}$ free and performing noisy inference updates. Figure 3 illustrates MCPC’s ability to learn non-linear multimodal distributions using the iris dataset (Pedregosa et al., 2011) and shows generative samples for MNIST. When comparing MCPC to a VAE on MNIST, both models produced samples of similar quality despite the VAE having twice the number of parameters. MCPC achieved a lower FID score (MCPC: $2.53^{\pm 0.17}$ vs. VAE: $4.19^{\pm 0.38}$), whereas the VAE attained a higher inception score (VAE: $7.91^{\pm 0.03}$ vs. MCPC: $7.13^{\pm 0.10}$).
In the associative memory (AM) experiments, we test how well the model is able to reconstruct an image already present in the training set, after it is provided with an incomplete or corrupted version of it, as done in a previous work (Salvatori et al., 2021). Fig. 4 show the results obtained by a PCN with 2 hidden layers of 512 neurons given noise or mask corrupted images. In Tab. 3, we study the memory capacity as the number of hidden layers increases. No visual difference between the recall and original images can be observed for MSE up to $0.005$. To evaluate efficiency we then trained a PCN with 5 hidden layers of 512 neurons on 500 TinyImagenet samples, with a batch size of 50 and 50 inference iterations during training. Training takes $0.40\pm 0.005$ seconds per epoch on an Nvidia V100 GPU.
@@ Discussion.
The results show that PC is able to perform generative tasks, as well as and associative memory ones using decoder-only architectures. Via inference, PCNs are able to encode complex probability distributions in their latent state which can be used to perform a variety of different tasks, as we have shown. Thus, compared to artificial neural networks, PCNs are more flexible and require only half the parameters to achieve similar performance. This comes at a higher computational cost due to the number of inference steps to perform. Future work should look into this issue, aiming at reducing the inference time by propagating the information more efficiently through the network.
@@ 5 Analysis and metrics
In this section, we report several metrics that we believe are important to understand the current state and challenges of training networks with PC and compare them with standard models trained with gradient-descent and backprop when suitable. More in detail, we discuss how regularly the energy flows into the model, and how stable training is when changing parameters, initializations, and optimizers. A better understanding of such phenomena would allow us to solve the current problems of PCNs and, hence, scale up to the training of larger models on more complex datasets.
5.1 Energy and stability
The first study we perform regards the initialization of the network states $h$, and how this influences the performance of the model. In the literature, they have been either initialized to be equal to zero, randomly initialized via a Gaussian prior (Whittington and Bogacz, 2017), and be initialized via a forward pass. This last technique has been the preferred option in machine learning papers as it sets the errors $\mathcal{\epsilon}_{l\neq L}=0$ at every internal layer of the model. This allows the prediction error to be concentrated in the output layer only, and hence be equivalent to the SE. To provide a comparison among the three methods, we have trained a $3$-layer feedforward model on FashionMNIST. The results, plotted in Fig. 5(a), show that forward initialization is indeed the better method, although the gap in performance shrinks the more iterations $T$ are performed.
Figure 5: (a) Highest test accuracy reported for different initialization methods and iteration steps $T$ used during training. (b) Energies per layer during inference of the best performing model (which has $\gamma=0.003$). (c) Decay in accuracy when increasing the learning rate of the states $\gamma$, tested using both SGD and Adam. (d) Imbalance between energies in the layers. All figures are obtained using a three layer model trained on FashionMNIST.
@@ 5.1 Energy and stability
Concentrating the total error of the model, and hence its energy, to the last layer as done when performing forward initialization, makes it hard for the model to then propagate such an energy back to the first layers. As reported in Fig. 5(b), we observe that the energy in the last layer is orders of magnitude larger than the one in the input layer, even after performing several inference steps. However, this behavior raises the question whether better initialization or optimization techniques could result in a more balanced energy distribution and thus better weight updates, as learning in this unbalanced energy regime has been shown problematic for more complex models (Alonso et al., 2024). An easy way of quickly propagating the energy through the network is to use learning rates equal to $1.0$ for the updates of the states. However, both the results reported in Fig. 5(c), as well as our large experimental analysis of Section 4 show that the best performance was consistently achieved for state learning rates $\gamma$ significantly smaller than $1.0$.
We hypothesize that the current training setup for PCNs favors small state learning rates that are sub-optimal to scale to deeper architectures. Fig. 5(d) shows the energy ratios for different state learning rates: when $\gamma\ll 1.0$, the ratios of energies between layers are small, $\nicefrac{{\epsilon_{l+1}^{2}}}{{\epsilon_{l}^{2}}}\ll 1$. The energy in the the first hidden layer is on average 6 orders of magnitude below the last layer for $\gamma=0.01$. While models trained with large $\gamma$ values achieve better energy propagation, they achieve lower accuracy as shown in Fig 5(c). Note that the decay in performance as function of increasing $\gamma$ is stronger for Adam despite being the overall better optimizer in our experiments. This suggests limitations in the current training techniques and possible direction for future improvements aimed at reducing the energy imbalance between layers. We provide implementation details and results on other datasets in Appendix D.
@@
We observed a further link between the weight optimizer and the structure of a PCN that might hinder the scalability of PC, that is, the influence of the hidden dimension on the performance of the model. To better study this, we trained feedforward PCNs with different hidden dimensions, state learning rates $\gamma$ and optimizers, and reported the results in Fig. 6. The results show that, when using Adam, the width strongly affects the values of the learning rate $\gamma$ for which the training process is stable. Interestingly, this phenomenon does not appear when using both the SGD optimizer, nor on standard networks trained with backprop. This behavioral difference with BP is unexpected and suggests the need for better optimization strategies for PCNs, as Adam was still the best choice in our experiments, but could be a bottleneck for larger architectures.
5.2 Properties of predictive coding networks
With PCX, it is straightforward to inspect and analyze several properties of PCNs. Here, we use $\mathcal{F}$ to differentiate between in-distribution (ID) and out-of-distribution (OOD) due to a semantic distribution shift Liu et al. (2020), as well as to compute the likelihood of a datasets Grathwohl et al. (2020). This can occur when samples are drawn from different, unseen classes, such as FashionMNIST samples under an MNIST setup Hendrycks and Gimpel (2017). To understand the confidence of the predictions of a PCN, we compare the distribution of the probability $p(x,\hat{y};\theta)$, for ID and OOD samples against these to the distribution of softmax values of the PCN classifier, and compute their negative log-likelihoods (NLL), according to the assumptions stated in Section 3, that is,
$\mathcal{F}=-\ln{p(x,y;\theta)}\implies p(x,y;\theta)=e^{-\mathcal{F}}.$ (4)
Our results in Fig. 7(a) demonstrate that a trained PCN classifier can effectively (1) assess OOD samples out-of-the-box, without requiring specific training for that purpose Yang et al. (2021), and (2) produce energy scores for ID and OOD samples that initially correlate with softmax values prior to the optimization of the states variables, $h$. However, after optimizing the states for $T$ inference steps, the scores for ID and OOD samples become decorrelated, especially for samples with lower softmax values as shown in Fig. 7(b). To corroborate this observation, we also present ROC curves for the most challenging samples, including only the lowest $25\%$ of the scores. As shown in Fig.7(c), the probability (i.e., energy-based) scores provide a more reliable assessment of whether samples are OOD. Experiment details and results on other datasets are provided in in Appendix E.
@@ 5.2 Properties of predictive coding networks
6 Computational Resources and Implementations Details
PCX is developed on top on JAX, focusing on performance and versatility. We plan to further improve its efficiency and expand its capabilities to support new developments in the field of backprop-free training techniques that align with the PC core principles. In relation to current alternatives, however, PCX is a very competitive choice. We measured the wall-clock time of our PCNs implementation against another existing open-source library (Song, 2024) used in many PC works (Song et al., 2024; Salvatori et al., 2021, 2022; Tang et al., 2023), as well as comparing it with equivalent BP-trained networks (developed also with PCX for a fair comparison). Tab. 4 reports the measured time per epoch, averaged over 5 trials, using a A100 GPU. Despite being a sub-optimal architecture in term of classification performance, we also tested on AlexNet (Krizhevsky et al., 2012), showing that our library can efficiently train models of more than 100 million parameters (having AlexNet $\approx 160$ million parameters). We also outperform alternative methods such as Eqprop: using the same architecture on CIFAR100, the authors report that one epochs takes $\approx 110$ seconds, while we take $\approx 5.5$ on the same hardware (Scellier et al., 2024). However, we stress that this is not an apple-to-apple comparison, as the authors are more concerned with simulations on analog circuits, rather than achieving optimal GPU usage.

### v2

@@ 4.1 Discriminative Mode
VGG-9
CIFAR-10 $78.12^{\pm 0.14}$ $75.33^{\pm 0.25}$ $76.90^{\pm 0.18}$ $85.90^{\pm 0.14}$ $87.19^{\pm 0.41}$ $79.02^{\pm 0.21}$ $89.18^{\pm 0.08}$ $\mathbf{90.02^{\pm 0.18}}$
CIFAR-100 (Top-1) $58.25^{\pm 0.13}$ $39.57^{\pm 0.18}$ $43.21^{\pm 0.21}$ $60.74^{\pm 0.75}$ $58.92^{\pm 1.61}$ $44.76^{\pm 0.40}$ $60.63^{\pm 0.28}$ $\mathbf{65.51^{\pm 0.23}}$
CIFAR-100 (Top-5) $83.28^{\pm 0.06}$ $66.90^{\pm 0.26}$ $71.13^{\pm 0.23}$ $83.19^{\pm 0.38}$ $81.56^{\pm 0.63}$ $72.88^{\pm 0.29}$ $\mathbf{85.25^{\pm 0.11}}$ $84.70^{\pm 0.28}$
Tiny ImageNet (Top-1) $39.64^{\pm 0.17}$ $21.78^{\pm 0.15}$ $23.62^{\pm 0.23}$ $41.59^{\pm 0.27}$ $31.5^{\pm 0.70}$ $26.34^{\pm 0.03}$ $\mathbf{45.66^{\pm 0.09}}$ $45.51^{\pm 0.15}$
Tiny ImageNet (Top-5) $64.60^{\pm 0.09}$ $44.43^{\pm 0.09}$ $46.89^{\pm 0.11}$ $66.15^{\pm 0.32}$ $54.67^{\pm 0.68}$ $50.48^{\pm 0.05}$ $\mathbf{69.65^{\pm 0.09}}$ $65.62^{\pm 0.17}$
ResNet-18
CIFAR-10 $43.19^{\pm 0.61}$ $53.74^{\pm 0.43}$ $62.45^{\pm 0.52}$ $62.33^{\pm 0.93}$ $55.29^{\pm 1.65}$ $70.44^{\pm 0.81}$ $92.83^{\pm 0.18}$ $\mathbf{93.21^{\pm 0.07}}$
CIFAR-100 (Top-1) $16.01^{\pm 0.42}$ $22.83^{\pm 0.38}$ $25.86^{\pm 0.86}$ $26.91^{\pm 0.55}$ $15.45^{\pm 1.7}$ $29.45^{\pm 1.36}$ $\mathbf{72.32^{\pm 0.26}}$ $71.89^{\pm 0.16}$
CIFAR-100 (Top-5) $40.67^{\pm 0.70}$ $50.18^{\pm 0.52}$ $53.80^{\pm 1.13}$ $55.57^{\pm 0.80}$ $39.42^{\pm 2.8}$ $56.70^{\pm 1.73}$ $\mathbf{92.14^{\pm 0.12}}$ $87.80^{\pm 0.18}$
Tiny ImageNet (Top-1) $09.52^{\pm 0.32}$ $14.19^{\pm 0.25}$ $15.79^{\pm 1.10}$ $15.95^{\pm 0.27}$ $04.40^{\pm 0.49}$ $06.19^{\pm 1.09}$ $\mathbf{58.00^{\pm 0.23}}$ $55.30^{\pm 0.16}$
Tiny ImageNet (Top-5) $26.21^{\pm 0.50}$ $34.55^{\pm 0.20}$ $37.36^{\pm 1.57}$ $37.76^{\pm 0.52}$ $14.30^{\pm 1.92}$ $16.51^{\pm 3.09}$ $\mathbf{79.94^{\pm 0.06}}$ $74.98^{\pm 0.36}$
We test the performance of PCNs on image classification tasks by comparing PC against BP, using both Squared Error (SE) and Cross Entropy (CE) loss, by adapting the energy function as described in Pinchetti et al. (2022). For the experiments on MNIST and FashionMNIST, we use feedforward models with $3$ hidden layers of $128$ hidden neurons, while for CIFAR10/100 and Tiny ImageNET, we compare ResNets and VGG-like models (He et al., 2016; Simonyan & Zisserman, 2014).
Results.
Table 1 shows that the best performing algorithms, at least on the most complex tasks, are the nudging ones (PN, NN, and CN). Among them, CN is almost always the best performing one, a result that is in line with previous findings in the Eqprop literature (Scellier et al., 2024). The only case where nudging algorithms are outperformed is on Tiny Imagenet on VGG7, where PC-CE performs better than them. However, the results obtained by PC-CE are still worse than the ones obtained by CN on VGG5. The recently proposed iPC, on the other hand, performs well on small architectures, as it is the best performing one on MNIST and FashionMNIST, but its performance worsens when it comes to the training of large architectures. More broadly, the performance of models of depth up to $7$ is comparable to those of backprop, while those of deeper models lag behind.
Discussion on depth.
An interesting observation is that all the best results for PC have been achieved using a VGG5, with the performance trend being VGG5 > VGG7 > VGG9 > ResNet, as shown in Fig 2. Conversely, we observe the opposite for backprop-trained models, with deeper models like VGG9 outperforming VGG5. A similar trend was observed in ResNet18 experiments, where PCNs yielded significantly lower test accuracies, with none of the models coming close to the performance of a VGG5. In contrast, backprop-trained ResNet18 models outperformed all previously tested VGG models, further emphasizing the gap in scalability between the two. Future work should investigate the reason of such a phenomenon, as scaling up to more complex datasets will require the use of much deeper architectures. In Section 5, we analyze possible causes, as well as comparing the wall-clock time of the different algorithms.
Figure 2: Test accuracies of different PC algorithms on the CIFAR10 dataset, using models of different depths.
@@ Results.
Figure 3: CIFAR10 image reconstruction via autoencoding convolutional networks. In order: original, PC, iPC, and BP.
Figure 4: Generative samples obtained by MCPC. Left: Contour plot of learned generative distribution compared to Iris data samples (x). Right: Samples obtained for a PCN. In order: unconditional generation, conditional generation (odd), conditional generation (even).
@@ 4.2 Generative Mode
In this section, we test the performance of PCNs on image generation tasks. We perform three different kinds of experiments: (1) generation from a posterior distribution; (2) generation via sampling from the learned joint distribution; and (3) associative memory retrieval. In the first case, we provide a test image $y$ to a trained model, run inference to compute a compressed representation $\bar{x}$ (stored in the latent vector $h_{0}$ at convergence), and produce a reconstructed $\bar{y}=h_{L}$ by performing a forward pass with $h_{0}=\bar{x}$). The models we consider have three layers, and we compare against autoencoders with a three-layer encoder/decoder structure (so, six layers in total). In the case of MNIST and FashionMNSIT we use feedforward layers, in the case of CIFAR10 and CelebA (de-)convolutional ones. The results in Tab. 2 and Fig. 3 report comparable performance, with a small advantage for PC compared to BP on the more complex tasks. In this case, iPC is the best performing algorithm, probably due to the small size of the considered models which allows for better stability.
Table 3: MSE (×$\times 10^{-4}$) of associative memory tasks given noisy (left) or masked (right) inputs as keys. Columns indicate the number of hidden neurons while rows shows the training images to memorize. Results over 5 seeds.
@@
Then, we tested the capability of PCNs to learn, and sample from, a complex probability distribution. MCPC extends PC by incorporating Gaussian noise to the activity updates of each neuron. This change enables a PCN to learn and generate samples analogous to a variational autoencoder (VAE). This change shifts the inference of PCNs from a variational approximation to Monte Carlo sampling of the posterior using Langevin dynamics. Data samples can be generated from the learned joint $P_{\theta}(h)$ by leaving all states $h_{l}$ free and performing noisy inference updates. Figure 4 illustrates MCPC’s ability to learn multimodal distributions using the iris dataset (Pedregosa et al., 2011) and shows generative samples for MNIST. When comparing MCPC to a VAE, both models produced samples of similar quality. MCPC achieved a lower FID score (MCPC: $2.53^{\pm 0.17}$ vs. VAE: $4.19^{\pm 0.38}$), whereas the VAE attained a higher inception score (VAE: $7.91^{\pm 0.03}$ vs. MCPC: $7.13^{\pm 0.10}$).
@@
In the associative memory (AM) experiments, we test how well the model is able to reconstruct a training image, after it is provided with an incomplete or corrupted version of it, as done in a previous work (Salvatori et al., 2021). Fig. 5 show the results obtained by a PCN with 2 hidden layers of 512 neurons given noise or mask corrupted images. In Tab. 3, we study the memory capacity as the number of hidden layers increases. No visual difference between the recall and original images can be observed for MSE up to $0.005$. To evaluate efficiency we then trained a PCN with 5 hidden layers of 512 neurons on 500 TinyImagenet samples, with a batch size of 50 and 50 inference iterations during training. Training takes $0.40\pm 0.005$ seconds per epoch on an Nvidia V100 GPU.
@@ Discussion.
The results show that PC is able to perform generative tasks, as well as associative memory ones using decoder-only architectures. Via inference, PCNs are able to encode complex probability distributions in their latent state which can be used to perform a variety of different tasks, as we have shown. While this highlights the flexibility of PCNs when used in the generative mode, this comes at a higher computational cost due to the number of inference steps to perform.
@@ 5 Analysis and metrics
In this section, we report several metrics that we believe are important to understand the current state and challenges of training networks with PC and compare them with standard models trained with gradient descent and backprop when suitable. The first study we perform analyzes how the initialization of the network states $h$ influences the performance of the model. In the literature, they have been either initialized to be equal to zero, randomly initialized via a Gaussian prior (Whittington & Bogacz, 2017), or initialized via a forward pass. This last technique has been the preferred option in machine learning papers as it sets the errors $\mathcal{\epsilon}_{l\neq L}=0$ at every internal layer of the model. This allows the prediction error to be concentrated in the output layer only, and hence be equivalent to the SE. To provide a comparison among the three methods, we have trained a $3$-layer feedforward model on FashionMNIST. The results, plotted in Fig. 6(a), show that forward initialization is indeed the better method, although the gap in performance shrinks the more iterations $T$ are performed.
Figure 6: (a): Highest test accuracy reported for different initialization methods and iteration steps $T$ used during training; (b): Energies per layer during inference of the best performing model (which has $\gamma=0.003$); (c) Decay in accuracy when increasing the learning rate of the states $\gamma$, tested using both SGD and Adam; (d) Imbalance between energies in the layers. Figures are obtained using a three layer model on FashionMNIST.
@@
Concentrating the total error of the model to the last layer makes it hard for the inference process to then propagate such an energy back to the first layers. As reported in Fig. 6(b), we observe that the energy in the last layer is orders of magnitude larger than the one in the input layer, even after performing several inference steps. An easy way of quickly propagating the energy through the network would be to use learning rates equal to $1.0$ for the updates of the states, that do not produce any energy imbalance, as also shown in Fig. 6(d). However, both the results reported in Fig. 6(b), as well as our large experimental analysis of Section 4 show that the best performance was consistently achieved for state learning rates $\gamma$ significantly smaller than $1.0$. This raises the question of whether better initialization or optimization techniques could result in a more balanced energy distribution and thus better weight updates.
To better understand how the energy propagation relates to the performance of the model, we have analyzed both the test accuracy and the ratio of the energies of subsequent layers as a function of the state learning rates $\gamma$. The results, reported in Fig 6(c,d), show that small learning rates lead to better performance, but also to large energy imbalances among layers. On the one hand, the energy in the first hidden layer is similar to that of the last layer for $\gamma=1$, and about $6$ orders of magnitude lower for $\gamma=0.01$. On the other hand, models trained with a learning rate of $\gamma=1$ achieve much worse performance. Such results show that the current training setup favors large energy imbalances among different layers, a problem that leads to exponentially small gradients when the depth of the model increases. We provide implementation details and results on other datasets in Appendix D.
@@
We have observed a link between the weight optimizer and the influence of the hidden dimension on the performance of the model. To better study this, we trained feedforward PCNs with different hidden dimensions, state learning rates $\gamma$ and optimizers, and reported the results in Fig. 7. The results show that, when using Adam, the width strongly affects the values of the learning rate $\gamma$ for which the training process is stable. Interestingly, this phenomenon does not appear when using both the SGD optimizer, nor on standard networks trained with backprop. This behavioral difference with BP is unexpected and suggests the need for better optimization strategies for PCNs, as AdamW was still the best choice in our experiments, but could be a bottleneck for larger architectures.
@@
6 Library, Resources and Implementations Details
In this section, we discuss PCX, the tool that we have used to perform the experiments, and that we release open source. PCX is developed on top of JAX, focusing on performance and versatility, and is built upon the following concepts: compatibility, modularity, and efficiency.
Compatibility. PCX shares the same philosophy of equinox (Kidger & Garcia, 2021), according to which models are just PyTrees. Consequently, it is fully compatible, using a complete functional approach, with both libraries and many other tools developed for JAX, such as diffrax (Kidger, 2021) and optax (DeepMind et al., 2020). To this end, it will be straightforward to implement novel development in deep learning into PCX. However, it also offers an imperative object-oriented interface, which allows researchers to build PCNs following a PyTorch-like style.
Modularity. Thanks to the object-oriented abstraction, we built the modular primitives that can be combined to create a PCN, mainly: a module class, representing abstract energy-based models; the vectorised nodes storing the states $h$; the optimizers, to perform the inference and learning process in a predictive coding network; and various standard Layers. Each benchmark we showcase in this work can be obtained by combining and configuring different blocks as needed.
Efficiency. PCX extensively relies on just-in-time compilation. From our initial benchmarks, we observed a speed-up of up to 50x when compiling a PCN. We believe that this stark difference is due to the nature of PC, which relies on multiple smaller operations compared to backpropagation, i.e., the $T$ inference step performed in each layer, and thus is more affected by the function calls overhead present in eager execution mode.
PCX offers a unified interface to test multiple variations of PC on several tasks. Our modular code base can easily be expanded in the future to support new variations of PC, as we show complete compatibility with existing variations and training techniques. This is different from, for example, the monolithic or low-level approaches used in (Song, 2024) and (Ororbia & Kifer, 2022), respectively.
6.1 Computational resources and limitations.
We measured the wall-clock time of our PCNs implementation against another existing open-source library (Song, 2024) used in many PC works (Song et al., 2024; Salvatori et al., 2021; 2022; Tang et al., 2023), as well as comparing it with equivalent BP-trained networks (developed also with PCX for a fair comparison). Tab. 4 reports the measured time per epoch, averaged over 5 trials, using a A100 GPU. We also outperform alternative methods such as Eqprop: using the same architecture on CIFAR100, the authors report that one epoch takes $\approx 110$ seconds, while we take $\approx 5.5$ on the same hardware (Scellier et al., 2024). However, this is not an apple-to-apple comparison, as the authors are more concerned with simulations on analog circuits, rather than achieving optimal GPU usage.